OpenStudy (anonymous):

Alexei wants to hang a mirror in his boat but the mirror and frame must not have an area larger than 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The border must have equal width on each side.) I WILL AWARD MEDALS.

OpenStudy (amistre64):

(3+w) (5+w) = area of the whole thing, which needs to be 19.25 or less (3+w) (5+w) <= 19.25

OpenStudy (anonymous):

@amistre64 it gives you answer choices 2x2 + 8x − 19.25 = 0 3x2 + 5x − 19.25 = 0 4x2 + 16x − 4.25 = 0 4x2 + 16x + 15 = 0

OpenStudy (amistre64):

i dont see any of those as a valid option.

OpenStudy (amistre64):

i see i forgot a little something tho

OpenStudy (amistre64):

each width attributes twice to a side

OpenStudy (anonymous):

Oh, well which seems most valid, I don't have any other options.

OpenStudy (amistre64):

(3+2w) (5+2w) = 19.25 thats better

OpenStudy (amistre64):

starts with a 4 and ends with a 4.25

OpenStudy (amistre64):